VIBRATIONAL SPECTRA AND AB-INITIO CALCULATIONS FOR THE STUDY OF INTRAMOLECULAR VIBRATIONAL REDISTRIBUTION IN THE CH-CHROMOPHORE IN CHD_{2}I

\underline{VERONIKA HORKÁ}, MARTIN QUACK, MARTIN WILLEKE, Laboratorium für Physikalische Chemie, ETH Zürich, CH-8093, Zürich, Switzerland.

Intramolecular vibrational energy redistribution (IVR) on the femto- to picosecond time scale is a central primary process in unimolecular chemical reaction dynamics. There are two main approaches towards its understanding. One is the analysis of the highly resolved vibrational spectra of polyatomic molecules^{a}^{,}^{b} combined with ab initio calculations. Another approach consists in time resolved femtosecond (fs) pump-probe experiments, which we have recently carried out for a series of alkyliodides^{c} including CHD_{2}I. For the latter molecule we present here results for IVR processes within the CH-chromophore. The infrared absorption spectrum of CHD_{2}I from 500 to 12000 cm^{−1} has been recorded with a Bomem Fourier transform infrared spectrometer equipped with a White type multi reflection cell and resolutions of 0.10 and 0.15 cm^{−1}. The CH-chromophore bands dominate the vibrational overtone spectrum of CHD_{2}I and we have assigned 24 of them. Vibrational variational calculations were carried out for the overtone spectrum of the CH-chromophore based on a 3-dimensional ab initio potential energy and electric dipole moment surface in normal coordinates. The anharmonic resonance structure between the CH-stretching and bending modes was investigated via an effective Hamiltonian for both experimental and theoretical spectra. Finally, we have calculated the time dependent dynamics on the fs time scale^{d}.

aM. Quack, Annu. Rev. Phys. Chem., 41, 839 (1990); M. Quack, Chapter 27 in Femtosecond Chemistry (ed. by J. Manz and L. Woeste), Verlag Chemie, Weinheim, 781 (1995); M. Quack, Chimia, 55 (10), 753 (2001) and references therein.

bJ. Pochert, M. Quack, J. Stohner, and M. Willeke, J. Chem Phys., 113, 2719 (2000).

cV.N. Krylov, M.V. Nikitchenko, M. Quack, and G. Seyfang, Proc. SPIE 5337, 178 (2004); V.N Krylov, A. Kushnarenko, E.

Miloglyadov, M. Quack, and G. Seyfang, Proc. SPIE, (2007) in press.

dM. Quack, and J. Stohner, J. Phys. Chem., 97(48), 12574 (1993).